Problems

1. The vapor pressure of pure water is 3.17 kPa. What will be the vapor pressure of a solution containing 18 g. of H2SO4 and 150 g of water?


2. If 75 g. of Lithium Chloride (ionic compound) is dissolved in 150 grams of water, what will be the freezing point of the solution?


3. What is the percent (volume/volume) if 12 ml of glycerol is added to 50 mL of water?


4. 0.50 grams of sodium sulfite is dissolved to make 0.050 mL of solution. What is the molarity of this solution?


5. Describe the process you would follow to prepare 150 ml of a 2.5 M solution of HCl if you started with a 5.0 M stock solution.


6. A gas has a solubility in water of 21.9 g/L at 15°C and 555 kPa of pressure. What is the pressure of the gas, in atm., (at the same temperature) if the solubility of the gas is 24 mg/L? 


1
Expert's answer
2021-04-05T07:47:32-0400


1)

Vapour =(150÷18)=8.33


2)

T = imK

∆T = change in freezing point temperature = ?

i = van't Hoff factor = 1 for lithium chloride since it is a non electrolyte and does not dissociate or ionize

Moles of lithium chloride="\frac{75}{420.4}=0.19moles"


"Molarity = \frac{0.19}{0.15}=1.27m"


Temp (freezing point)= 1×1.27×1.86=2.3622



3)

"Volume/volume=\frac{12}{50}×100=24" %





4)


Moles of sodium sulfite"=\frac{0.50}{126}=3.97×10^{-3}"


Molarity of sodium sulfite "= \frac{0.050}{1000}(3.97×10^{-3})=1.985×10^{-7}"